OpenStudy (anonymous):

Find the difference. 4d2e - d2e

OpenStudy (anonymous):

that would be 3d2e

OpenStudy (anonymous):

because of alike terms we can subtract.

OpenStudy (anonymous):

okay thanks! but do you think you can explain what to do?

OpenStudy (anonymous):

yes, let us take take this for example 2x-x what would you do here?

OpenStudy (anonymous):

recall since the x doesn't have a

OpenStudy (anonymous):

a number the x is 1

OpenStudy (anonymous):

soo 2x-1x?

OpenStudy (anonymous):

so you have 1x

OpenStudy (anonymous):

opps that is 2x-1x= 1x

OpenStudy (anonymous):

so just subtract??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okayy! thanks.

OpenStudy (anonymous):

only if they have a like terms like x or y's something the same. don't try to this if there is x and z and y have you can not subtract

OpenStudy (anonymous):

6rs -2rs like terms?

OpenStudy (anonymous):

yes, because they both rs so therefor you can subtract

OpenStudy (anonymous):

so the answer should be 4rs

OpenStudy (anonymous):

i was boutta put that... lol

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

!:)

OpenStudy (anonymous):

any other questions

OpenStudy (anonymous):

4ab2 - 3ab2 =1?

OpenStudy (anonymous):

4ab2-3ab2=1ab2 not just one

OpenStudy (anonymous):

don't forget when you add y's or x the answer still has the terms in it.

OpenStudy (anonymous):

so because in this we have 4ab2-3ab2= 1ab2 not just 1 otherwise the answer iswrong

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thast not a option

OpenStudy (anonymous):

??

OpenStudy (anonymous):

1ab2 is not a answer

OpenStudy (anonymous):

opps my mistake it's just one in this case sorry..

OpenStudy (anonymous):

that is right because the term when subtracted go to zero so therefor it's just one

OpenStudy (anonymous):

so im right ?

OpenStudy (anonymous):

yes in this case , again sorry for the typo though

OpenStudy (anonymous):

its cool

OpenStudy (anonymous):

3x2y - x2y

OpenStudy (anonymous):

okay what do we know about this problem? we have 3x2y-x2y= these are a like terms you tell me

OpenStudy (anonymous):

and no the terms aren't there still , that was my mistake there..

OpenStudy (anonymous):

so it should be 2

OpenStudy (anonymous):

you get it?

OpenStudy (anonymous):

yeah sorry! to busy on youtube @godorovg

OpenStudy (anonymous):

5wx2 - 2wx2= 3wx2 @godorovg ????/

OpenStudy (anonymous):

it should be 3